# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

### v2

@@ 3 Discussion
Courts understand this factor to concern the extent to which works produced through copying serve as market substitutes for the original author’s work (?, ?). Our study has shown that AI-generated excerpts from in-context prompted models pre-trained on vast internet corpora (including millions of copyrighted works) were strongly disfavored by expert readers for both quality and stylistic fidelity, though lay readers showed no clear preference. By contrast, when models are fine-tuned on curated datasets—consisting solely of individual author’s complete works—both experts and lay readers decisively preferred the AI-generated excerpts over human-written examples. Moreover, the fine-tuned excerpts proved almost undetectable as AI-generated text (particularly when compared to outputs from in-context prompted models), while consistently surpassing human writing in blind evaluations.
@@
The Copyright Office has recognized that such predicate copying may cause cognizable market harm through competing works that the inputs enable, potentially flooding the market and causing “market dilution” (?). While acknowledging this “market dilution” approach to the fourth fair use factor as “uncharted territory,” the Office determined that both the statutory language and underlying concerns of the Copyright Act warrant this inquiry.999“…The speed and scale at which AI systems generate content pose a serious risk of diluting markets for works of the same kind as in their training data. That means more competition for sales of an author’s works and more difficulty for audiences in finding them. If thousands of AI-generated romance novels are put on the market, fewer of the human-authored romance novels that the AI was trained on are likely to be sold…Market harm can also stem from AI models’ generation of material stylistically similar to works in their training data…” The Office emphasized that the effect of the copying impacts extant works by putting them in competition with AI-generated outputs. Copyrighted works become fodder for new productions targeting the same markets. Crucially, the Office did not claim that competing AI outputs copy the inputted works; rather, it examined the economic consequences of the predicate copying that enables these competing outputs. This focus on inputs remains essential because, absent the initial copying, no infringement action exists against flooding markets with independently generated works—if a thousand humans write romance novels after reading Barbara Cartland’s novels, they compete but do not infringe. The Copyright Office’s expansive interpretation of “potential market for or value of the copied work” suggests that fair use might not excuse predicate copying even when it doesn’t show up in the end product, if the copying’s effect substitutes for source works.
@@
In Kadrey v. Meta (?), an infringement action brought by 13 book authors against the copying of their books into the database underpinning Meta’s Llama LLM, Judge Chhabria granted summary judgment to Meta but accepted the theory of market dilution:101010[I]ndirect substitution is still substitution: If someone bought a romance novel written by an LLM instead of a romance novel written by a human author, the LLM-generated novel is substituting for the human-written one…This case involves a technology that can generate literally millions of secondary works, with a miniscule fraction of the time and creativity used to create the original works it was trained on. No other use…has anything near the potential to flood the market with competing works the way that LLM training does. Judge Chhabria effectively provided a road map for what authors would have to show to persuade a court that the AI inputs diluted their markets: First, is the AI system “capable of generating” substitutional books? Second, what are the markets for the plaintiffs’ books, and do the AI-generated books compete in those markets? “Third, what impact does this competition actually have on sales of the books it competes with?…Whatever the effects have been thus far, are they likely to increase in the future, as more and more AI-generated books are written, and as LLMs get better and better at writing human-like text? Fourth, how does the threat to the market for the plaintiffs’ books in a world where LLM developers can copy those books compare to the threat to the market for the plaintiffs’ books in a world where the developers can’t copy them?”
@@
Our study’s findings concerning reader preferences between human-authored and AI-generated works bear on all four considerations. They also demonstrate how LLMs have already gotten “better and better at writing human-like text.” While Judge Chhabria speculated that the distinctiveness of an author’s style renders works by well-known authors less susceptible to substitution, [^11] Judge Chhabria observed that market dilution would vary by author prominence: established authors with dedicated readerships (like Agatha Christie) would likely face minimal substitution, while AI-generated books could crowd out lesser-known or emerging authors, potentially preventing “the next Agatha Christie from getting noticed or selling enough books to keep writing.” our work suggests otherwise. If readers in fact prefer AI-generated emulations of authors whose market value lies in their distinctive voices, then the prospect of competition, especially from outputs of fine-tuned datasets, appears to be considerable. The comparatively low production costs of AI-generated texts relative to paying human authors (as shown in Figure 4C) further enhances the likelihood that AI platforms will in fact dilute the market for human-authored work.
@@
A reasonable solution might allow the inclusion of copyrighted works in the general-purpose dataset, but would require the model to implement guardrails that would disable it from generating non-parodic imitations of individual authors’ oeuvres (?, ?, ?). [^12] As examples of guardrails, AI developers have implemented “refusal protocols” blocking outputs when prompts request content “in the style of” specific authors. Further, current reinforcement learning techniques can be easily modified to steer models away from stylistic imitation. Another solution, particularly where the lower quality of in-context prompting reduces the prospect of market dilution, might be to condition a ruling of fair use on the prominent disclosure of the output’s AI origin. This solution assumes that the public, informed that the output was not human-authored, will be less inclined to select the AI substitute; transparency should diminish the competition between human-authored and machine-generated offerings. The solution also assumes that in-context prompting will not in the future produce outputs that readers will prefer to human-authored text. Improvements in LLMs (and/or increases in the number of works copied into training data) may come to belie that assumption.
@@ 4 Methods
We recruited 28 candidates from top MFA programs (Iowa Writers’ Workshop, Helen Zell Writers’ Program at the University of Michigan, MFA Program in Creative Writing at New York University, Columbia University School of the Arts), paying each \$75 for writing a single excerpt. These MFA candidates and LLMs emulated the style/voice of 50 award-winning authors representing diverse cultural backgrounds and distinct literary voices (full list in Table 2 in SI). The writing prompt provided to MFA candidates contained (i) 20 sample excerpts spanning an author’s complete body of work (ii) textual descriptions of the author’s distinctive style and voice (iii) detailed content specifications about the original author written excerpt to be emulated. The selection of the 50 authors and all the writing prompts were developed in collaboration with five English Literature PhD students who analyzed each author’s literary voice and created the verbalized style descriptions (prompt in Figure 5 in SI).
@@
These <<Human-AI>> pairs for both conditions were evaluated by 28 experts (the same MFA candidates) and 131 lay readers recruited from Prolific, [^13] https://www.prolific.com/ one of the leading crowdsourcing platforms for research participants. Experts never evaluated their own excerpts. Each pair was assessed by three experts and five lay readers, with majority voting determining final judgments. Inter-rater agreement was quantified using Fleiss’ kappa. In total, we obtained 2,400 pairwise evaluations (1,200 quality, 1,200 style) for AI Condition 1 and 1,440 evaluations (720 quality, 720 style) for AI Condition 2. For quality evaluation, we showed the <<Human-AI>> pair alone; for style evaluation, we included the original author written excerpt alongside the pair (see Figures 9 and 10 in SI). Both experts and lay readers also provided 2-3 sentence explanations grounded in textual evidence to justify their choices (?) (see Figures 1720 in SI).
To ensure annotation quality, we implemented attention checks including timestamp recording to prevent rushing. Additionally, we screened responses using Pangram, [^14] https://www.pangram.com/ a state-of-the-art AI detection tool, and excluded participants who used generative AI in their responses (?). Our study was approved by the University of Michigan IRB (HUM00264127) and was preregistered at OSF. [^15] https://osf.io/zt4ad Informed consent was obtained from all participants.
We tested hypotheses H1 (baseline LLMs vs. human writers) and H2 (fine-tuned GPT-4o vs. human writers) using logistic regression with CR2 cluster-robust standard errors clustered by reader. For H1, we compared human writing to the average performance across GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro in the in-context prompting condition. For H2, we directly compared fine-tuned GPT-4o to human writing. Contrasts were computed separately for expert and lay readers within each outcome (style, quality), with Holm correction applied across reader-group contrasts within each hypothesis-outcome combination. For H3 (AI detection and preference), we modeled the relationship between Pangram AI-detection scores and preference, testing whether fine-tuning attenuated detection-based penalties via setting $\times$ detection score interactions. All the model specifications are described in detail in SI.
Limitations
Our recruitment was mostly restricted to American creative writing programs and further study needs to be done across creative writing programs outside the US. While our pre-dedicated pool of 50 writers consisted of some writers who do not write in English, our experiments on style/voice emulation were done based on their English translation. Creative writing often depends on intrinsic motivation. While we offered MFA students a lucrative rate for writing the excerpts, it’s unclear if monetary incentives actually enhanced their creative output, since intrinsic motivation typically drives the best artistic work. Last but not least our experiments were conducted at a shorter excerpt level and conclusions cannot be drawn for long form text. In it’s current form AI is unable to generate long form text that’s thematically coherent unlike humans. While we foresee a situation where human can collaborate with a finetuned AI model to create competing long form works, experimental evidence is required to make any broader claims.
Code and Data availability
@@
All analysis and figure-generation code alongside human data is available upon request. Core analyses can be reproduced by running the numbered R scripts in sequence. Analyses were conducted in R 4.3.1 with key packages including clubSandwich and emmeans. Exact package versions and run instructions are provided in SI, Section S10.
@@ Acknowledgements
Author contributions
TC, PSD: Conceptualization; Methodology; Data Analysis; Writing (original draft); Writing (review and editing).
JCG: Writing (original draft); Writing (review and editing).
Competing interests
The authors declare no competing interests.
Materials & correspondence
Correspondence should be addressed to Tuhin Chakrabarty (tchakrabarty@cs.stonybrook.edu), Jane C. Ginsburg (ginsburg@law.columbia.edu), or Paramveer Dhillon (dhillonp@umich.edu).
@@
Table 2 shows the list of 50 authors that were chosen by English Literature Ph.D. students. These chosen author list consists of canon-plus-global giants (Ernest Hemingway, Virginia Woolf, William Faulkner, Gabriel Garcia Márquez, Stephen King, Haruki Murakami, Kazuo Ishiguro), a strong set of contemporary prominent literary voices (Margaret Atwood, Ian McEwan, Jonathan Franzen, Colson Whitehead, George Saunders, Louise Erdrich, Octavia Butler, Salman Rushdie, Maya Angelou, Percival Everett), and a group of critically acclaimed / emerging authors (Ottessa Moshfegh, Tony Tulathimutte, Roxane Gay ). Additionally our author list is culturally diverse where several of the authors write primarily in a non English language (Han Kang, Yoko Ogawa, Annie Ernaux). Roughly two‑thirds (34) have secured at least one major international or national prize (e.g., Nobel, Booker, Pulitzer, National Book Award, MacArthur Fellowship, Women’s Prize, International Booker, Hugo/Nebula). 8 of the authors are Nobel Prize winners in Literature and 8 are Pulitzer Prize winners.
@@ 18 Ian McEwan (✓) 36 Percival (✓)Everett
Larger context windows in LLMs models allow for processing significantly more information at once, leading to improved accuracy, understanding, and complex reasoning capabilities. Taking advantage of this feature in GPT4-o, Claude-3.5-Sonnet and Gemini-1.5-Pro, we design long context prompt that first demonstrates 20 sample excerpts written by a given author, followed by their style/voice verbalized in text and finally the content of the excerpt to be emulated (See Figure 5). The same prompt was provided to both experts( MFA candidates) and LLMs.
@@
Figure 6: The pipeline used to finetune ChatGPT on an authors entire oeuvre
For finetuning we bought digital ePub versions of these authors’ books and transformed them into plain text files.If the epub file was basically a wrapper around scanned page images, then we ignored that epub.The number of books written by each author vary a lot. For instance Tony Tulathimutte has written only two books Private Citizens and Rejection so we could only finetune GPT4-o on two of them. While for Haruki Murakami we could finetune on 22 books A Wild Sheep Chase, After Dark, After the Quake, Blind Willow, Sleeping Woman, Colorless Tsukuru Tazaki and his Years of Pilgrimage, Dance Dance Dance, First Person Singular, Hard-boiled Wonderland and the End of the World, Hear the Wind Sing, Kafka on the Shore, Killing Commendatore, Men Without Women, Norwegian Wood, Novelist as a Vocation, One and Two, Pinball, 1973, South of the Border, West of the Sun, Sputnik Sweetheart, The City and Its Uncertain Walls, The Elephant Vanishes, The Wind-Up Bird Chronicle, What I Talk About When I Talk About Running, Wind/Pinball: Two Novels
Finetuning on books isn’t a straightforward process. Each book is typically 50,000 to 80,000 words long with some exceptions. Using such long sequences wastes capacity because models still struggle with very long inputs and long‑context adaptation is non‑trivial; To preserve general capabilities in Supervised Finetuning it’s generally advised to have diverse, shorter samples. Breaking a book into context‑independent excerpts increases batch diversity and the number of distinct gradient signals per token budget. At the same time it also reduces overfitting to a single narrative flow while still injecting the salient stylistic/local patterns.
Our entire finetuning pipeline can be seen in Figure 6. We first convert the epub files to txt using https://github.com/kevinboone/epub2txt2. We then segment the entire book into context independent excerpts. At a first pass we naively split the entire book text by existing double-newlines and rejoin them to enforce excerpt size bounds(250-650 words). For the rare cases where the naive splitting would lead to excerpts longer than 650 words we used GPT4o to segment them again using the prompt Segment it into excerpts of minimum length 300-350 words such that each excerpt is grammatical from the start and doesn’t feel abruptly cut off. There should be zero deletion and break into excerpts at grammatically natural places. Maintain the original word count. Avoid breaking into too many small excerpts. Start directly. Don’t say Here’s or Here is ….
After obtaining context excerpts we extract content details from them using GPT4-o by prompting it Describe in detail what is happening in this excerpt. Mention the characters and whether the voice is in first or third person for majority of the excerpt.Maintain the order of sentences while describing.. Figure 7 shows a sample paragraph from Shame written by Annie Ernaux and the extracted content. Once we obtain the extracted content we finetune GPT4-o using their fine-tuning API with the following input-output pair Write a [[n]] word excerpt about the content below emulating the style and voice of [[authorname]]: [[excerpt]] as seen in Figure 6. This technique is commonly referred to as instruction back-translation (?)
@@
After finetuning is completed at inference time we can then simply generate an excerpt conditioned on a similar instruction that contains a novel content. We specifically excluded the original author written excerpt used for our experiments during fine-tuning so that the model does not get an unfair advantage. We also ensured that the generated outputs do not contain any memorized snippets from the original author written excerpt. In the rare occasion that a model regurgitated verbatim snippets/ngrams from original author written text we resampled it and manually verified that there is no verbatim overlap before using it for evaluation. While supervised finetuning in most cases ensures that the generated output contains all the information in the extracted content details, in the rare occasion that it does not we resample/regenerate it again. Last but not least sometimes supervised finetuning can lead to ungrammatical output or output with minor inconsistencies. To make sure these mistakes don’t impact human evaluation, we performed a post processing step using GPT4-o using the following prompt Fix grammar, tense, typo, spelling or punctuation error or any other awkward construction/ logical inconsistency
@@
For Margaret Atwood, Kazuo Ishiguro, Salman Rushdie and Haruki Murakami we fine-tuned GPT4-o for 1 epoch as they had multiple books/longer books (For instances . For the rest of the authors we fine-tuned for 3 epochs. We use default finetuning parameters set by OpenAI in addition to setting batch size as 1 and LR multiplier as 2. Table 16 shows that finetune models don’t regurgitate a lot of verbatim text from the training corpus (author’s entire oeuvre). We should also note that some of these words are also part of the Content in the instruction which would anyway appear in the text and should not be penalized. For all generation(prompting or fine-tuning default temperature=1.0 was used)
@@
S2.1 Expert and Lay Recruitment Demographic
All experts(MFA candidates) recruited were currently residing in USA. This was a requirement for payment purposes as depending on how much an individual makes as a part of the study requires them to declare it as taxable income. Paying experts who do not have a ITIN or SSN would be challenging for logistical reasons. 65% of our expert did not identify as cis gender men. In terms of ethnicities our experts identified as South Asian, White, East Asian and Black. For lay readers recruited from Prolific we restricted ourselves to English speaking countries (only USA and UK). We also required participants to be born there, have a 100% acceptance rate and for everyone to be college educated.
S2.2 Evaluation Interface
@@
Figures 9 and 10 show the evaluation interface shown to both lay and expert readers. Readers read two excerpts for quality evaluation and three excerpts for stylistic fidelity evaluation given that the stylistic emulation is evaluated with respect to the original author written excerpt.
@@
Without observing why a choice was made (no rationale/side information), the latent factors driving that preference are often unclear. Such hidden user context, demographic/value variation, task ambiguity, or near‑tie similarity—remain inflates effective label noise. Providing rationales or auxiliary side information has been empirically shown to improve label reliability, data efficiency, and debiasing (?). This led us to ask for reasons that support the preference choice. For lay readers, reasons also act as proxy for understanding if the reader is doing the task sincerely and not choosing preference randomly.
S2.3 How much is faithful style/voice emulation important for good writing?

### v3

@@ 2 Discussion
Courts understand this factor to concern the extent to which works produced through copying serve as market substitutes for the original author’s work (?, ?). Our study has shown that AI-generated excerpts from in-context prompted models pre-trained on vast internet corpora (including millions of copyrighted works) were strongly disfavored by expert evaluators for both quality and stylistic fidelity, though lay readers showed no clear preference. By contrast, when models are fine-tuned on curated datasets—consisting solely of individual author’s complete works—both experts and lay readers decisively preferred the AI-generated excerpts over human-written examples. Moreover, the fine-tuned excerpts proved almost undetectable as AI-generated text (particularly when compared to outputs from in-context prompted models), while consistently surpassing human writing in blind evaluations.
@@
The U.S. Copyright Office has recognized that such predicate copying may cause cognizable market harm through competing works that the inputs enable, potentially flooding the market and causing “market dilution” (?). While acknowledging this “market dilution” approach to the fourth fair use factor as “uncharted territory,” the Office determined that both the statutory language and underlying concerns of the Copyright Act warrant this inquiry. “…The speed and scale at which AI systems generate content pose a serious risk of diluting markets for works of the same kind as in their training data. That means more competition for sales of an author’s works and more difficulty for audiences in finding them. If thousands of AI-generated romance novels are put on the market, fewer of the human-authored romance novels that the AI was trained on are likely to be sold…Market harm can also stem from AI models’ generation of material stylistically similar to works in their training data…” The Office emphasized that the effect of the copying impacts extant works by putting them in competition with AI-generated outputs. Copyrighted works become fodder for new productions targeting the same markets. Crucially, the Office did not claim that competing AI outputs copy the inputted works; rather, it examined the economic consequences of the predicate copying that enables these competing outputs. This focus on inputs remains essential because, absent the initial copying, no infringement action exists against flooding markets with independently generated works—if a thousand humans write romance novels after reading Barbara Cartland’s novels, they compete but do not infringe. The Copyright Office’s expansive interpretation of “potential market for or value of the copied work” suggests that fair use might not excuse predicate copying even when it doesn’t show up in the end product, if the copying’s effect substitutes for source works.
@@
In Kadrey v. Meta (?), an infringement action brought by 13 book authors against the copying of their books into the database underpinning Meta’s Llama LLM, Judge Chhabria granted summary judgment to Meta but accepted the theory of market dilution: “[I]ndirect substitution is still substitution: If someone bought a romance novel written by an LLM instead of a romance novel written by a human author, the LLM-generated novel is substituting for the human-written one…This case involves a technology that can generate literally millions of secondary works, with a miniscule fraction of the time and creativity used to create the original works it was trained on. No other use…has anything near the potential to flood the market with competing works the way that LLM training does.” Judge Chhabria effectively provided a road map for what authors would have to show to persuade a court that the AI inputs diluted their markets: First, is the AI system capable of generating such substitutional books now or in the near future? Second, what are the markets for the plaintiffs’ books, and do the AI-generated books compete in those markets? Third, what impact does any such competition have on sales—does it displace those works entirely or merely chip away at demand, and are the effects likely to grow as AI-generated books proliferate and models improve? Fourth, how does the threat to the market for the plaintiffs’ books in a world where LLM developers can copy those books compare to the threat to the market for the plaintiffs’ books in a world where the developers can’t copy them?
@@
Our study’s findings concerning reader preferences between human-authored and AI-generated works bear on all four considerations. They also demonstrate how LLMs have already gotten “better and better at writing human-like text.” While Judge Chhabria speculated that the distinctiveness of an author’s style renders works by well-known authors less susceptible to substitution, he observed that market dilution would vary by author prominence. Established authors with dedicated readerships (like Agatha Christie) would likely face minimal substitution, while AI-generated books could crowd out lesser-known or emerging authors, potentially preventing “the next Agatha Christie from getting noticed or selling enough books to keep writing”. Our work suggests otherwise. If readers in fact prefer AI-generated emulations of authors whose market value lies in their distinctive voices, then the prospect of competition, especially from outputs of fine-tuned datasets, appears to be considerable. The comparatively low production costs of AI-generated texts relative to paying human authors (as shown in Figure 4C) further enhances the likelihood that AI platforms will in fact dilute the market for human-authored work.
@@
A reasonable solution might allow the inclusion of copyrighted works in the general-purpose dataset, but would require the model to implement guardrails that would disable it from generating non-parodic imitations of individual authors’ oeuvres (?, ?, ?).As examples of guardrails, AI developers have implemented “refusal protocols” blocking outputs when prompts request content “in the style of” specific authors. Further, current reinforcement learning techniques can be easily modified to steer models away from stylistic imitation. Another solution, particularly where the lower quality of in-context prompting reduces the prospect of market dilution, might be to condition a ruling of fair use on the prominent disclosure of the output’s AI origin. This solution assumes that the public, informed that the output was not human-authored, will be less inclined to select the AI substitute; transparency should diminish the competition between human-authored and machine-generated offerings. The solution also assumes that in-context prompting will not in the future produce outputs that readers will prefer to human-authored text. Improvements in LLMs (and/or increases in the number of works copied into training data) may come to belie that assumption.
@@ 3 Methods
3.1 Experimental Setup
We recruited 28 candidates from top MFA programs in the United States, including from programs such as Iowa Writers Workshop, Helen Zell Writers’ Program at University of Michigan, MFA Program in Creative Writing at New York University, Columbia University School of Arts, and MFA in Creative Writing at BU, paying them \$75 for writing each excerpt. These MFA candidates, along with the LLMs described earlier, emulated the style and voice of 50 award-winning authors representing diverse cultural backgrounds and distinct literary voices (full list in Table 2 in SI). The writing prompt provided to MFA candidates contained three components: (i) 20 sample excerpts spanning an author’s complete body of work, (ii) textual descriptions of the author’s distinctive style and voice, and (iii) detailed content specifications about the original author-written excerpt to be emulated. The selection of the 50 authors and the development of all writing prompts were completed in collaboration with five English Literature PhD students who analyzed each author’s literary voice and created the verbalized style descriptions (prompt shown in Figure 5 in SI).
3.1.1 In-context Prompting
@@
3.1.2 Fine-tuning
For AI Condition 2 (fine-tuning), we selected 30 living authors from our pool. This decision was made specifically to examine the potential economic impact of generative AI on the livelihoods of living authors while also considering the substantial computational costs associated with fine-tuning models on individual authors. Fine-tuning on books presents several technical challenges. Each book typically contains 50,000 to 80,000 words, and using such long sequences as training examples wastes computational capacity because models still struggle with very long inputs and long-context adaptation remains non-trivial. To preserve general capabilities in supervised fine-tuning, it is generally advised to use diverse, shorter samples. Breaking books into context-independent excerpts increases batch diversity and the number of distinct gradient signals per token budget, while also reducing overfitting to a single narrative flow and still capturing the salient stylistic and local patterns characteristic of each author’s voice.
[…]
3.2 Human Participants for Writing and Evaluation Task
[…]
Our expert sample represents the emerging literary elite whose professional judgment shapes contemporary publishing. Among our recruited MFA candidates, several have since been awarded Stanford Stegner Fellowships (one of the most competitive fiction writing fellowships in the United States), received Rhodes Scholarships, assumed editorial positions at prestigious literary magazines such as Joyland, and secured publishing contracts with major houses including W.W. Norton and Harper Collins. These accomplishments underscore that our experts represent not merely MFA candidates, but established and emerging voices whose aesthetic judgments carry professional authority in the literary marketplace.
All recruited experts were residing in the United States at the time of the study. This geographic restriction was necessary for administrative compliance, as U.S. tax regulations require declaration of compensation above certain thresholds, and the university’s payment infrastructure could not accommodate international participants without Individual Taxpayer Identification Numbers (ITINs) or Social Security Numbers (SSNs). Despite this constraint, our expert sample demonstrated substantial demographic diversity: 65% did not identify as cisgender men, and participants self-identified across multiple ethnic backgrounds, including South Asian, White, East Asian, and Black. This diversity helps ensure that our findings are not artifacts of a narrow demographic perspective.
For the 131 lay readers recruited from Prolific, we restricted participation to native-born residents of English-speaking countries (USA and UK only) ensuring that all evaluators had native-level fluency in English. Additional eligibility criteria included maintaining a 100% task acceptance rate on the platform (indicating high performance on previous tasks) and having a college degree (ensuring a baseline level of literacy). We compensated the lay readers according to Prolific’s wage standards with approximately \$15 per hour for their evaluation time. Unlike experts, lay readers represented general educated audiences without specialized training in creative writing or literary analysis; hence their compensation was lower than experts.
3.3 Blind Evaluation
[…]
Beyond selecting a preferred excerpt, both experts and lay readers provided 2–3 sentence explanations grounded in textual evidence to justify their choices  (?) (see Figures 2023 in SI for examples). We required these written rationales for two reasons. First, without observing why a choice was made, the latent factors driving that preference remain unclear—hidden user context, demographic or value variation, task ambiguity, or near-tie similarity can inflate effective label noise. Providing rationales or auxiliary side information has been empirically shown to improve label reliability, data efficiency, and debiasing  (?). Second, for lay readers specifically, written explanations serve as a quality control mechanism, helping us identify participants who may be selecting preferences randomly rather than engaging seriously with the task. Readers who provided generic or off-topic rationales, or whose explanations contradicted their stated preferences, were flagged for potential exclusion.
To ensure annotation quality, we implemented several attention checks. We recorded timestamps for each evaluation to identify and exclude participants who rushed through tasks (spending less than 30 seconds per comparison). Additionally, we screened all written rationales using Pangram, [^5] https://www.pangram.com/ a state-of-the-art AI detection tool, and excluded any participants who used generative AI to compose their responses  (?). This screening was necessary because AI-generated rationales would undermine the validity of human preference judgments. Our study was approved by the University of Michigan IRB (HUM00264127) and was preregistered at OSF. [^6] https://osf.io/zt4ad Informed consent was obtained from all participants. As noted earlier, we compensated experts \$75 per 10-excerpt batch for quality judgments and \$100 per 10-excerpt batch for stylistic fidelity evaluations (the higher rate reflecting the additional cognitive load of comparing against original author excerpts).
3.4 Empirical Analyses
We tested three preregistered hypotheses using logistic regression models with CR2 cluster-robust standard errors clustered by reader. Clustering accounts for the fact that each reader evaluated multiple excerpt pairs, which induces within-reader correlation in preference judgments. The CR2 adjustment provides more accurate inference in settings with a modest number of clusters (28 expert readers, 131 lay readers)  (?).
For H1 (can baseline LLMs match human expert performance?), we compared human writing to the average performance across GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro in the in-context prompting condition. Because each human excerpt was paired with only one AI excerpt from a single model, we pooled AI excerpts across the three models and tested whether readers showed systematic preference for human versus AI writing on average. For H2 (can fine-tuned models match human expert performance?), we directly compared fine-tuned GPT-4o excerpts to human writing. Contrasts were computed separately for expert and lay readers within each outcome (stylistic fidelity, writing quality), yielding four tests per hypothesis (2 reader groups $\times$ 2 outcomes). To control the family-wise error rate across these multiple comparisons, we applied Holm correction separately within each hypothesis-outcome combination. This procedure adjusts $p$-values to maintain valid inference when testing multiple contrasts that address the same research question.
For H3 (does AI detectability correlate with human preference, and does fine-tuning remove this correlation?), we modeled the relationship between Pangram AI-detection scores and reader preference. Specifically, we tested whether higher detection scores (indicating more AI-like text) predicted lower preference for AI excerpts in the in-context prompting condition, and whether fine-tuning attenuated this penalty via AI condition $\times$ detection score interactions. A significant negative coefficient on detection score in the in-context condition, coupled with a significant positive interaction term, would indicate that fine-tuning removes the detectability-based preference penalty that AI excerpts suffer under in-context prompting. All model specifications, including the full regression equations, covariate adjustments, and robustness checks, are described in detail in SI Sections S4–S8.
Data availability
Anonymized trial-level data supporting the findings of this study will be made publicly available on paper acceptance. If you need them for research purpose, please email the first author. Additional details are provided in SI, Section S10.
Code availability
@@ 3.4 Empirical Analyses
All analysis and figure-generation code will be made publicly available on paper acceptance. Core analyses can be reproduced by running the numbered R scripts in sequence. Analyses were conducted in R 4.3.1 with key packages including clubSandwich and emmeans. Exact package versions and run instructions are provided in SI, Section S10.
@@ Acknowledgements
Table 2 shows the list of 50 authors that were chosen by English Literature Ph.D. students. These chosen author list consists of canon-plus-global giants (Ernest Hemingway, Virginia Woolf, William Faulkner, Gabriel Garcia Márquez, Stephen King, Haruki Murakami, Kazuo Ishiguro), a strong set of contemporary prominent literary voices (Margaret Atwood, Ian McEwan, Jonathan Franzen, Colson Whitehead, George Saunders, Louise Erdrich, Octavia Butler, Salman Rushdie, Maya Angelou, Percival Everett), and a group of critically acclaimed / emerging authors (Ottessa Mosfegh, Tony Tulathimutte, Roxane Gay). Additionally our author list is culturally diverse where several of the authors write primarily in a non-English language (Han Kang, Yoko Ogawa, Annie Ernaux). Roughly two‑thirds (34) have secured at least one major international or national prize (e.g., Nobel, Booker, Pulitzer, National Book Award, MacArthur Fellowship, Women’s Prize, International Booker, Hugo/Nebula). 8 of the authors are Nobel Prize winners in Literature and 8 are Pulitzer Prize winners.
@@ 18 Ian McEwan (✓) 36 Percival (✓)Everett
Our writing prompt can be seen in Figure 5). The same prompt was provided to both experts( MFA candidates) and LLMs.
@@
Figure 6: The pipeline used to finetune ChatGPT on an author’s entire oeuvre
For finetuning we bought digital ePub versions of these authors’ books and transformed them into plain text files. If the epub file was basically a wrapper around scanned page images, then we ignored that epub. The number of books written by each author vary a lot. For instance Tony Tulathimutte has written only two books Private Citizens and Rejection so we could only finetune GPT-4o on two of them. While for Haruki Murakami we could finetune on 22 books A Wild Sheep Chase, After Dark, After the Quake, Blind Willow, Sleeping Woman, Colorless Tsukuru Tazaki and his Years of Pilgrimage, Dance Dance Dance, First Person Singular, Hard-boiled Wonderland and the End of the World, Hear the Wind Sing, Kafka on the Shore, Killing Commendatore, Men Without Women, Norwegian Wood, Novelist as a Vocation, One and Two, Pinball, 1973, South of the Border, West of the Sun, Sputnik Sweetheart, The City and Its Uncertain Walls, The Elephant Vanishes, The Wind-Up Bird Chronicle, What I Talk About When I Talk About Running, Wind/Pinball: Two Novels
Our entire finetuning pipeline can be seen in Figure 6. We first convert the epub files to txt using https://github.com/kevinboone/epub2txt2. We then segment the entire book into context independent excerpts. At a first pass we naively split the entire book text by existing double-newlines and rejoin them to enforce excerpt size bounds(250-650 words). For the rare cases where the naive splitting would lead to excerpts longer than 650 words we used GPT-4o to segment them again using the prompt Segment it into excerpts of minimum length 300-350 words such that each excerpt is grammatical from the start and doesn’t feel abruptly cut off. There should be zero deletion and break into excerpts at grammatically natural places. Maintain the original word count. Avoid breaking into too many small excerpts. Start directly. Don’t say Here’s or Here is ….
@@
After finetuning is completed at inference time we can then simply generate an excerpt conditioned on a similar instruction that contains a novel content. We specifically excluded the original author written excerpt used for our experiments during fine-tuning so that the model does not get an unfair advantage. We also ensured that the generated outputs do not contain any memorized snippets from the original author written excerpt. In the rare occasion that a model regurgitated verbatim snippets/ngrams from original author written text we resampled it and manually verified that there is no verbatim overlap before using it for evaluation. While supervised fine-tuning in most cases ensures that the generated output contains all the information in the extracted content details, in the rare occasion that it does not we resample/regenerate it again. Last but not least sometimes supervised fine-tuning can lead to ungrammatical output or output with minor inconsistencies. To make sure these mistakes don’t impact human evaluation, we performed a post processing step using GPT-4o using the following prompt Fix grammar, tense, typo, spelling or punctuation error or any other awkward construction/ logical inconsistency
@@
For Margaret Atwood, Kazuo Ishiguro, Salman Rushdie and Haruki Murakami we finetuned GPT-4o for 1 epoch as they had multiple books/longer books. For the rest of the authors we finetuned for 3 epochs. We use default finetuning parameters set by OpenAI in addition to setting batch size as 1 and LR multiplier as 2. Table 19 shows that finetune models don’t regurgitate a lot of verbatim text from the training corpus (author’s entire oeuvre). We should also note that some of these words are also part of the Content in the instruction which would anyway appear in the text and should not be penalized. For all generation(prompting or fine-tuning default temperature=1.0 was used)
S1.4 Why recruit MFAs as expert writers
As mentioned earlier MFA is typically recognized as a terminal degree for practitioners of visual art, design, dance, photography, theatre, film/video, new media, and creative writing—meaning that it is considered the highest degree in its field, qualifying an individual to become a professor at the university level in these disciplines. As can be seen in Figures 11, 9 and 11 most top writers have an MFA degree. While there are exceptional writers without any formal training the space for such individuals is enormous and difficult to systematically sample, whereas MFA graduates can be more readily identified and tracked through institutional records.
Figure 9: Table showing how several award winning authors from US all had MFA degree
Figure 10: Table showing a few accomplished alumni with MFA degree from programs where we recruited our participants
Figure 11: Table showing breakdown of New Yorkers 20 under 40 Fiction List
@@
S2.1 Evaluation Interface
@@
Figures 12 and 13 shows the evaluation interface shown to both lay and expert readers. Readers read two excerpts for quality evaluation and three excerpts for stylistic fidelity evaluation given that the stylistic emulation is evaluated with respect to the original author written excerpt.
@@
S2.2 How much is faithful style/voice emulation important for good writing?